OpenStudy (anonymous):

Find the linear approximation of the function f(x) = square root of 1 − x at a = 0.

OpenStudy (anonymous):

L(x) = f(a) + f '(a)(x-a) f(x) = sqrt(1- x) f(a) = f(0) = 1 f '(x) = (1/2)(-1 / sqrt(1-x)) f '(a) = f '(0) = ?